# On the Mode-Seeking Properties of Langevin Dynamics

### v2

@@ A.1 Proof of Theorem 1: Langevin Dynamics under Gaussian Mixtures
Therefore we complete the proof of Proposition 3. ∎
Finally, combining Propositions 2 and 3 finishes the proof of Theorem 1.
A.2 Proof of Corollary 1
By the definition of total variation distance, we have
$\text{TV}(\hat{P}_{t},P)=\sup_{A}|\hat{P}_{t}(A)-P(A)|.$
Specifically, by choosing the event $A$ as $\left\{{{\mathbf{x}}:\forall i\in[k],\,\left\|{{\mathbf{x}}-{\bm{\mu}}_{i}}\right\|^{2}\geq\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d}\right\}$, from Theorem 1 we know $\hat{P}_{t}(A)\geq 1-T\cdot\exp(-\Omega(d))$. On the other hand, by Lemma 1 we have
$P(A)$ $=\sum_{i=0}^{k}w_{i}P^{(i)}(A)$
$\leq w_{0}+\sum_{i=0}^{k}w_{i}\exp\left(-\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{\max}^{2}}\right)d\right)$
$=w_{0}+(1-w_{0})\exp\left(-\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{\max}^{2}}\right)d\right).$
Combining the two bounds, we obtain a lower bound on the total variation distance
$\text{TV}(\hat{P}_{t},P)\geq\hat{P}_{t}(A)-P(A)\geq(1-w_{0})\left(1-\frac{T}{\exp(\Omega(d))}\right).$
A.3 Proof of Theorem 2: Langevin Dynamics under Sub-Gaussian Mixtures
@@
For constants ${\nu}_{0},{\nu}_{i},c_{{\nu}},c_{L}$ satisfying Assumptions 2.iii. and 2.iv., we have $\frac{(1-c_{{\nu}}){\nu}_{0}^{2}-{\nu}_{i}^{2}}{2(1-c_{{\nu}})}>0$ and $\log\frac{c_{{\nu}}{\nu}_{i}^{2}}{(c_{L}^{2}+c_{{\nu}}c_{L}){\nu}_{0}^{2}}-\frac{{\nu}_{i}^{2}}{2(1-c_{{\nu}}){\nu}_{0}^{2}}+\frac{(1-c_{{\nu}}){\nu}_{0}^{2}}{2{\nu}_{i}^{2}}>0$ are both positive constants.
@@
where ${\mathbf{R}}\in{\mathbb{R}}^{d\times r}$ an orthonormal basis of the vector space $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$ and ${\mathbf{N}}\in{\mathbb{R}}^{d\times n}$ an orthonormal basis of the null space of $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$. To show $\left\|{{\mathbf{x}}_{t}-{\bm{\mu}}_{i}}\right\|^{2}>\left(\frac{{\nu}_{0}^{2}}{2}+\frac{{\nu}_{\max}^{2}}{2(1-c_{{\nu}})}\right)d$, it suffices to prove $\left\|{\mathbf{n}_{t}}\right\|^{2}>\left(\frac{{\nu}_{0}^{2}}{2}+\frac{{\nu}_{\max}^{2}}{2(1-c_{{\nu}})}\right)d$. By Proposition 2, if ${\mathbf{x}}_{0}$ is initialized in the distribution $P^{(0)}$, i.e., ${\mathbf{x}}_{0}\sim P^{(0)}$, since ${\nu}_{0}^{2}>\frac{1}{1-c_{{\nu}}}{\nu}_{\max}^{2}$, with probability at least $1-\exp(-\Omega(d))$ we have
$\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\left(\frac{3{\nu}_{0}^{2}}{4}+\frac{{\nu}_{\max}^{2}}{4(1-c_{{\nu}})}\right)d.$ (13)
Then, conditioned on $\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\left(\frac{3{\nu}_{0}^{2}}{4}+\frac{{\nu}_{\max}^{2}}{4(1-c_{{\nu}})}\right)d$, the following proposition shows that $\left\|{\mathbf{n}_{t}}\right\|$ remains large with high probability.
Proposition 4.
Consider a distribution $P$ satisfying Assumption 2. We follow the Langevin dynamics for $T=\exp({\mathcal{O}}(d))$ steps. Suppose that the initial sample satisfies $\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\left(\frac{3{\nu}_{0}^{2}}{4}+\frac{{\nu}_{\max}^{2}}{4(1-c_{{\nu}})}\right)d$, then with probability at least $1-T\cdot\exp(-\Omega(d))$, we have that $\left\|{\mathbf{n}_{t}}\right\|^{2}>\left(\frac{{\nu}_{0}^{2}}{2}+\frac{{\nu}_{\max}^{2}}{2(1-c_{{\nu}})}\right)d$ for all $t\in\{0\}\cup[T]$.
Proof of Proposition 4.
@@
To show $\frac{\left\|{\nabla_{\mathbf{x}}P^{(i)}({\mathbf{x}})}\right\|}{P({\mathbf{x}})}\leq\exp(-\Omega(d))$, from Assumptions 2.ii. and 2.iii. we have
@@
Armed with the above lemmas, we are now ready to establish Proposition 4 by induction. Please note that we also apply some lemmas from the proof of Theorem 1 by substituting ${\nu}_{\max}^{2}$ with $\frac{{\nu}_{\max}^{2}}{1-c_{{\nu}}}$. Suppose the theorem holds for all $T$ values of $11,\cdots,T-$ 1. We consider the following 3 cases:
@@
Therefore we complete the proof of Proposition 4. ∎
Finally, combining equation 13 and Proposition 4 finishes the proof of Theorem 2.
Appendix B Theoretical Analysis on Annealed Langevin Dynamics
B.1 Annealed Langevin Dynamics in Gaussian Mixtures
In the following Theorem 3 we extend the result in Theorem 1 to annealed Langevin dynamics with bounded noise levels.
Theorem 3.
Consider a data distribution $P$ satisfying Assumption 1. We follow annealed Langevin dynamics for $T=\exp({\mathcal{O}}(d))$ steps with noise levels $c_{{\sigma}}\geq{\sigma}_{0}\geq\cdots\geq{\sigma}_{T}\geq 0$ for constant $c_{{\sigma}}>0$. In addition, assume for all $i\in[k]$, $\left\|{{\bm{\mu}}_{i}-{\bm{\mu}}_{0}}\right\|^{2}\leq\frac{{\nu}_{0}^{2}-{\nu}_{i}^{2}}{2}\left(\log\left(\frac{{\nu}_{i}^{2}+c_{{\sigma}}^{2}}{{\nu}_{0}^{2}+c_{{\sigma}}^{2}}\right)-\frac{{\nu}_{i}^{2}+c_{{\sigma}}^{2}}{2{\nu}_{0}^{2}+c_{{\sigma}}^{2}}+\frac{{\nu}_{0}^{2}+c_{{\sigma}}^{2}}{2{\nu}_{i}^{2}+c_{{\sigma}}^{2}}\right)d$. Suppose that the sample is initialized in $P^{(0)}_{{\sigma}_{0}}$, then with probability at least $1-T\cdot\exp(-\Omega(d))$, we have $\left\|{{\mathbf{x}}_{t}-{\bm{\mu}}_{i}}\right\|^{2}>\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}+2{\sigma}_{t}^{2}}{2}d$ for all $t\in\{0\}\cup[T]$ and $i\in[k]$.
Proof of Theorem 3.
Proposition 5.
Proof of Proposition 5.
By the definition in equation 2, we have
Similar to the proof of Theorem 1, we decompose
First, by Proposition 2, suppose that the sample is initialized in the distribution $P^{(0)}_{{\sigma}_{0}}$, then with probability at least $1-\exp(-\Omega(d))$, we have
$\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\frac{3({\nu}_{0}^{2}+{\sigma}_{0}^{2})+({\nu}_{\max}^{2}+{\sigma}_{0}^{2})}{4}d=\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}+4{\sigma}_{0}^{2}}{4}d.$ (20)
Then, with the assumption that the initialization satisfies $\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}+4{\sigma}_{0}^{2}}{4}d$, the following proposition similar to Proposition 3 shows that $\left\|{\mathbf{n}_{t}}\right\|$ remains large with high probability.
Proposition 6.
Consider a data distribution $P$ satisfies the constraints specified in Theorem 3. We follow annealed Langevin dynamics for $T=\exp({\mathcal{O}}(d))$ steps with noise level $c_{{\sigma}}\geq{\sigma}_{0}\geq{\sigma}_{1}\geq{\sigma}_{2}\geq\cdots\geq{\sigma}_{T}\geq 0$ for some constant $c_{{\sigma}}>0$. Suppose that the initial sample satisfies $\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}+4{\sigma}_{0}^{2}}{4}d$, then with probability at least $1-T\cdot\exp(-\Omega(d))$, we have that $\left\|{\mathbf{n}_{t}}\right\|^{2}>\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}+2{\sigma}_{t}^{2}}{2}d$ for all $t\in\{0\}\cup[T]$.
Proof of Proposition 6.
We prove Proposition 6 by induction. Suppose the theorem holds for all $T$ values of $11,\cdots,T-$ 1. We consider the following 3 cases:
  • If there exists some $t\in[T]$ such that $\delta_{t}>{\nu}_{0}^{2}+{\sigma}_{t}^{2}$, by Lemma 3 we know that with probability at least $1-\exp(-\Omega(d))$, we have $\left\|{\mathbf{n}_{t}}\right\|^{2}\geq\frac{3({\nu}_{0}^{2}+{\sigma}_{t}^{2})+({\nu}_{\max}^{2}+{\sigma}_{t}^{2})}{4}d=\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}+4{\sigma}_{t}^{2}}{4}d$, thus the problem reduces to the two sub-arrays $\mathbf{n}_{0},\cdots,\mathbf{n}_{t-1}$ and $\mathbf{n}_{t},\cdots,\mathbf{n}_{T}$, which can be solved by induction.
  • Suppose $\delta_{t}\leq{\nu}_{0}^{2}+{\sigma}_{t}^{2}$ for all $t\in[T]$. If there exists some $t\in[T]$ such that $\left\|{\mathbf{n}_{t-1}}\right\|^{2}>36({\nu}_{0}^{2}+{\sigma}_{t-1}^{2})d\geq 36({\nu}_{0}^{2}+{\sigma}_{t}^{2})d$, by Lemma 5 we know that with probability at least $1-\exp(-\Omega(d))$, we have $\left\|{\mathbf{n}_{t}}\right\|^{2}\geq({\nu}_{0}^{2}+{\sigma}_{t}^{2})d>\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}+4{\sigma}_{t}^{2}}{4}d$, thus the problem similarly reduces to the two sub-arrays $\mathbf{n}_{0},\cdots,\mathbf{n}_{t-1}$ and $\mathbf{n}_{t},\cdots,\mathbf{n}_{T}$, which can be solved by induction.
    Since ${\nu}_{0}>{\nu}_{i}$ and $c_{{\sigma}}\geq{\sigma}_{t}$ for all $t\in\left\{{0}\right\}\cup[T]$, we have $\frac{{\nu}_{i}^{2}+c_{{\sigma}}^{2}}{{\nu}_{0}^{2}+c_{{\sigma}}^{2}}\geq\frac{{\nu}_{i}^{2}+{\sigma}_{t}^{2}}{{\nu}_{0}^{2}+{\sigma}_{t}^{2}}$. Notice that for function $f(z)=\log z-\frac{z}{2}+\frac{1}{2z}$, we have $\frac{{\,\mathrm{d}}}{{\,\mathrm{d}}z}f(z)=\frac{1}{z}-\frac{1}{2}-\frac{1}{2z^{2}}=-\frac{1}{2}\left(\frac{1}{z}-1\right)^{2}\leq 0$. Thus, by the assumption
    $\left\|{{\bm{\mu}}_{i}-{\bm{\mu}}_{0}}\right\|^{2}\leq\frac{{\nu}_{0}^{2}-{\nu}_{i}^{2}}{2}\left(\log\left(\frac{{\nu}_{i}^{2}+c_{{\sigma}}^{2}}{{\nu}_{0}^{2}+c_{{\sigma}}^{2}}\right)-\frac{{\nu}_{i}^{2}+c_{{\sigma}}^{2}}{2{\nu}_{0}^{2}+c_{{\sigma}}^{2}}+\frac{{\nu}_{0}^{2}+c_{{\sigma}}^{2}}{2{\nu}_{i}^{2}+c_{{\sigma}}^{2}}\right)d,$
    we have that for all $t\in[T]$,
    $\left\|{{\bm{\mu}}_{i}-{\bm{\mu}}_{0}}\right\|^{2}\leq\frac{{\nu}_{0}^{2}-{\nu}_{i}^{2}}{2}\left(\log\left(\frac{{\nu}_{i}^{2}+{\sigma}_{t}^{2}}{{\nu}_{0}^{2}+{\sigma}_{t}^{2}}\right)-\frac{{\nu}_{i}^{2}+{\sigma}_{t}^{2}}{2{\nu}_{0}^{2}+{\sigma}_{t}^{2}}+\frac{{\nu}_{0}^{2}+{\sigma}_{t}^{2}}{2{\nu}_{i}^{2}+{\sigma}_{t}^{2}}\right)d.$
    Conditioned on $\left\|{\mathbf{n}_{t-1}}\right\|^{2}>\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}+2{\sigma}_{t-1}^{2}}{2}d$ for all $t\in[T]$, by Lemma 6 we have that for $T=\exp({\mathcal{O}}(d))$,
    By Lemma 8 we have that with probability at least $1-\exp(-\Omega(d))$,
    $\left\|{\hat{\mathbf{n}}_{T}}\right\|^{2}\geq\frac{5{\nu}_{0}^{2}+3{\nu}_{\max}^{2}+8{\sigma}_{T}^{2}}{8}d.$
    Combining the two inequalities implies the desired bound
    $\left\|{\mathbf{n}_{T}}\right\|\geq\left\|{\hat{\mathbf{n}}_{T}}\right\|-\left\|{\hat{\mathbf{n}}_{T}-\mathbf{n}_{T}}\right\|>\sqrt{\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}+2{\sigma}_{T}^{2}}{2}d}.$
    Hence by induction we obtain $\left\|{\mathbf{n}_{t}}\right\|^{2}>\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}+2{\sigma}_{t}^{2}}{2}d$ for all $t\in\left\{{0}\right\}\cup[T]$ with probability at least
Therefore we complete the proof of Proposition 6. ∎
Finally, combining equation 20 and Proposition 6 finishes the proof of Theorem 3. ∎
B.2 Annealed Langevin Dynamics in Sub-Gaussian Mixtures
Finally, we slightly modify Assumption 2 and extend our results to annealed Langevin dynamics (with bounded noise levels) under sub-Gaussian mixtures in Theorem 4.
Assumption 3.
  1. i.
  2. ii.
  3. iii.
  4. iv.
  5. v.
Theorem 4.
Consider a data distribution $P$ satisfying Assumption 3. We follow annealed Langevin dynamics for $T=\exp({\mathcal{O}}(d))$ steps with noise levels $c_{{\sigma}}\geq{\sigma}_{0}\geq\cdots\geq{\sigma}_{T}\geq 0$. Suppose the sample is initialized in $P^{(0)}_{{\sigma}_{0}}$, then with probability at least $1-T\cdot\exp(-{\mathcal{O}}(d))$, we have $\left\|{{\mathbf{x}}_{t}-{\bm{\mu}}_{i}}\right\|^{2}>\left(\frac{{\nu}_{0}^{2}+{\sigma}_{t}^{2}}{2}+\frac{{\nu}_{\max}^{2}+{\sigma}_{t}^{2}}{2(1-c_{{\nu}})}\right)d$ for all $t\in\{0\}\cup[T]$ and $i\in[k]$.

### v3

@@ A.2 Proof of Theorem 2
    $(1-(T-1)\exp(-d/300))\cdot(1-\exp(-d/300))\geq 1-T\exp(-d/300).$
Therefore we complete the proof of Theorem 2.
A.3 Proof of Theorem 3
Proof of Theorem 3.
Proposition 1.
Proof of Proposition 1.
By the definition in equation 2, we have
Similar to the proof of Theorem 2, we decompose
By Definition 2, since $\mathbf{n}_{t}$ is the projection onto the null space of $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$, we have
$\left\|{{\mathbf{x}}_{t}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}=\min_{\lambda_{1},\cdots,\lambda_{k}}\left\|{{\mathbf{x}}_{t}-\sum_{i=1}^{k}\lambda_{i}{\bm{\mu}}_{i}}\right\|=\left\|{\mathbf{n}_{t}}\right\|.$
We prove Theorem 3 by induction. Suppose the theorem holds for all $T$ values of $11,\cdots,T-$ 1. We consider the following 3 cases:
  • If there exists some $t\in[T]$ such that $\delta_{t}>{\nu}_{0}^{2}+{\sigma}_{t}^{2}$, by Lemma 9 we know that with probability at least $1-\exp\left(-\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8({\nu}_{0}^{2}+{\sigma}_{t}^{2})}\right)^{2}\frac{d}{2}\right)\geq 1-\exp(-d/32)$, we have $\left\|{\mathbf{n}_{t}}\right\|^{2}\geq\frac{3({\nu}_{0}^{2}+{\sigma}_{t}^{2})+({\nu}_{\max}^{2}+{\sigma}_{t}^{2})}{4}d=\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}+4{\sigma}_{t}^{2}}{4}d$, thus the problem reduces to the two sub-arrays $\mathbf{n}_{0},\cdots,\mathbf{n}_{t-1}$ and $\mathbf{n}_{t},\cdots,\mathbf{n}_{T}$, which can be solved by induction.
  • Suppose $\delta_{t}\leq{\nu}_{0}^{2}+{\sigma}_{t}^{2}$ for all $t\in[T]$. If there exists some $t\in[T]$ such that $\left\|{\mathbf{n}_{t-1}}\right\|^{2}>36({\nu}_{0}^{2}+{\sigma}_{t-1}^{2})d\geq 36({\nu}_{0}^{2}+{\sigma}_{t}^{2})d$, by Lemma 11 we know that with probability at least
    $1-\exp\left(-\left(\log\frac{{\nu}_{i}^{2}+{\sigma}_{t}^{2}}{{\nu}_{0}^{2}+{\sigma}_{t}^{2}}-\frac{{\nu}_{i}^{2}+{\sigma}_{t}^{2}}{2({\nu}_{0}^{2}+{\sigma}_{t}^{2})+\frac{{\nu}_{0}^{2}+{\sigma}_{t}^{2}}{2({\nu}_{i}^{2}+{\sigma}_{t}^{2})}}\right)\frac{d}{4}\right)-\frac{4}{\sqrt{2\pi d}}\exp\left(-\frac{d}{32}\right)$
    $\geq 1-\exp(-0.01d),$
    we have $\left\|{\mathbf{n}_{t}}\right\|^{2}\geq({\nu}_{0}^{2}+{\sigma}_{t}^{2})d>\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}+4{\sigma}_{t}^{2}}{4}d$, thus the problem similarly reduces to the two sub-arrays $\mathbf{n}_{0},\cdots,\mathbf{n}_{t-1}$ and $\mathbf{n}_{t},\cdots,\mathbf{n}_{T}$, which can be solved by induction.
    Conditioned on $\left\|{\mathbf{n}_{t-1}}\right\|^{2}>\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}+2{\sigma}_{t-1}^{2}}{2}d$ for all $t\in[T]$, by Lemma 12 we have that for $T\leq\exp(d/150)$,
    By Lemma 14 we have
    $\left\|{\hat{\mathbf{n}}_{T}}\right\|^{2}\geq\frac{5{\nu}_{0}^{2}+3{\nu}_{\max}^{2}+8{\sigma}_{T}^{2}}{8}d$
    with probability at least
    $1-\exp\left(-\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{0}^{2}+8{\sigma}_{0}^{2}}\right)^{2}\frac{d}{2}\right)-\frac{4\sqrt{7}}{\sqrt{\pi d}}\exp\left(-\frac{({\nu}_{0}^{2}-{\nu}_{\max}^{2})^{2}d}{32({\nu}_{0}^{2}+{\sigma}_{0}^{2})(3{\nu}_{0}^{2}+{\nu}_{\max}^{2}+4{\sigma}_{0}^{2})}\right)$
    $\geq 1-\exp\left(-\frac{d}{512}\right)-\frac{4\sqrt{7}}{\sqrt{\pi d}}\exp\left(-\frac{d}{448}\right)\geq 1-\exp\left(-\frac{d}{1500}\right).$
    Combining the two inequalities implies the desired bound
    $\left\|{\mathbf{n}_{T}}\right\|\geq\left\|{\hat{\mathbf{n}}_{T}}\right\|-\left\|{\hat{\mathbf{n}}_{T}-\mathbf{n}_{T}}\right\|>\sqrt{\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}+2{\sigma}_{T}^{2}}{2}d}\geq\sqrt{\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d}.$
    Hence by induction we obtain $\left\|{\mathbf{n}_{t}}\right\|^{2}>\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d$ for all $t\in\left\{{0}\right\}\cup[T]$ with probability at least
    $(1-(T-1)\exp(-d/1500))\cdot(1-\exp(-d/1500))\geq 1-T\exp(-d/1500).$
Therefore we complete the proof of Theorem 3. ∎
Appendix B Iteration Complexity of Langevin Dynamics in sub-Gaussian Mixtures
A probability distribution $p({\mathbf{z}})$ of dimension $d$ is defined as a sub-Gaussian distribution with parameter ${\nu}^{2}$ if, given the mean vector ${\bm{\mu}}:=\mathbb{E}_{{\mathbf{z}}\sim p}[{\mathbf{z}}]$, the moment generating function (MGF) of $p$ satisfies the following inequality for every vector ${\bm{\alpha}}\in{\mathbb{R}}^{d}$:
Assumption 1.
  1. i.
  2. ii.
    $P^{(i)}$ is differentiable and $\nabla P^{(i)}({\bm{\mu}}_{i})=\bm{0}_{d}$,
  3. iii.
  4. iv.
  5. v.
The feasibility of Assumption 1.v. is validated by Lemma 15 in Appendix B.1. With Assumption 1, we show the hardness of Langevin dynamics under sub-Gaussian distributions in Theorem 5 and defer the proof to Appendix B.1.
Theorem 5.
Consider a data distribution $P$ satisfying Assumption 1. We initialize the sample ${\mathbf{x}}_{0}$ such that $\left\|{{\mathbf{x}}_{0}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}^{2}\geq\left(\frac{3{\nu}_{0}^{2}}{4}+\frac{{\nu}_{\max}^{2}}{4(1-c_{{\nu}})}\right)d$ and apply Langevin dynamics for $T$ steps, then
${\mathbb{P}}\left(\left\|{{\mathbf{x}}_{T}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}^{2}\geq\left(\frac{{\nu}_{0}^{2}}{2}+\frac{{\nu}_{\max}^{2}}{2(1-c_{{\nu}})}\right)d\right)\geq 1-T\cdot\exp\left(-\Omega(d)\right).$
Then, we slightly modify Assumption 1 and extend our results to annealed Langevin dynamics (with bounded noise levels) under sub-Gaussian mixtures in Theorem 6. The proof of Theorem 6 is deferred to Appendix B.2.
Assumption 2.
  1. i.
  2. ii.
  3. iii.
  4. iv.
  5. v.
Theorem 6.
Consider a data distribution $P$ satisfying Assumption 2. We initialize the sample ${\mathbf{x}}_{0}$ such that $\left\|{{\mathbf{x}}_{0}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}^{2}\geq\left(\frac{3{\nu}_{0}^{2}+3c_{{\sigma}}^{2}}{4}+\frac{{\nu}_{\max}^{2}+c_{{\sigma}}^{2}}{4(1-c_{{\nu}})}\right)d$ and apply annealed Langevin dynamics for $T$ steps with noise levels $c_{{\sigma}}\geq{\sigma}_{0}\geq\cdots\geq{\sigma}_{T}\geq 0$, then
${\mathbb{P}}\left(\left\|{{\mathbf{x}}_{T}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}^{2}\geq\left(\frac{{\nu}_{0}^{2}}{2}+\frac{{\nu}_{\max}^{2}}{2(1-c_{{\nu}})}\right)d\right)\geq 1-T\cdot\exp\left(-\Omega(d)\right).$
We noticed that a central requirement of Theorems 2 and 5 is that the initial sample ${\mathbf{x}}_{0}$ must be far from the low-variance modes. In the following Theorem 7, we relax this constraint by considering low-variance modes $P^{(1)},P^{(2)},\cdots,P^{(k)}$ with random mean vectors, as characterized by Assumption 3. The proof of Theorem 7 is deferred to Appendix B.3
Assumption 3.
Consider a data distribution $P:=\sum_{i=0}^{k}w_{i}P^{(i)}$, where $k\geq 1$ and $w_{i}>0$ are positive constants such that $\sum_{i=0}^{k}w_{i}=1$. Suppose the density of mode 0 is lower bounded by $P^{(0)}({\mathbf{x}})\geq(2\pi{\nu}_{0}^{2})^{-d/2}\exp\left(-\frac{(1+c_{0})\left\|{{\mathbf{x}}}\right\|^{2}}{2{\nu}_{0}^{2}}\right)$ for some constant ${\nu}_{0}$ and $c_{0}\geq 0$. In addition, assume $\log P^{(0)}({\mathbf{x}})$ is concave and $\left\|{\nabla_{\mathbf{x}}\log P^{(0)}(\bm{0}_{d})}\right\|=\exp(o(d))$, and its score function $\nabla_{\mathbf{x}}\log P^{(0)}({\mathbf{x}})$ is $L_{0}$-Lipschitz. For all $i\in[k]$, suppose $P^{(i)}$ satisfies
  1. i.
    the mean ${\bm{\mu}}_{i}$ of $P^{(i)}$ is i.i.d. uniform over an $\ell_{2}$ ball ${\mathbb{S}}$ centered at $\bm{0}_{d}$ of radius $r$,
  2. ii.
    $P^{(i)}$ is a sub-Gaussian distribution with parameter ${\nu}_{i}^{2}$.
  3. iii.
    $P^{(i)}$ is differentiable and $\nabla P^{(i)}({\bm{\mu}}_{i})=\bm{0}_{d}$,
  4. iv.
  5. v.
    ${\nu}_{i}$ satisfies $\left(\frac{1-c_{{\nu}}}{2{\nu}_{i}^{2}}-\frac{1+c_{0}}{{\nu}_{0}^{2}}\right)\left(\frac{{\nu}_{0}^{2}}{2}+\frac{{\nu}_{i}^{2}}{2(1-c_{{\nu}})}\right)-\frac{1+c_{0}}{{\nu}_{0}^{2}}r^{2}-\frac{1}{2}\log\frac{c_{{\nu}}{\nu}_{i}^{2}}{(c_{L}^{2}+c_{{\nu}}c_{L}){\nu}_{0}^{2}}>0$ for some constant $c_{{\nu}}\in(0,1)$
Theorem 7.
Consider a data distribution $P$ satisfying Assumption 3. For any initial sample ${\mathbf{x}}_{0}$, we follow Langevin dynamics for $T$ steps with step size $\delta_{t}\leq 4/L_{0}$, then
${\mathbb{P}}\left(\left\|{{\mathbf{x}}_{T}}\right\|^{2}\geq\left(\frac{{\nu}_{0}^{2}}{2}+\frac{{\nu}_{\max}^{2}}{2(1-c_{{\nu}})}\right)d\right)\geq 1-T\cdot\exp\left(-\Omega(d)\right).$
B.1 Proof of Theorem 5
Proof of Theorem 5.
@@ B.1 Proof of Theorem 5
For constants ${\nu}_{0},{\nu}_{i},c_{{\nu}},c_{L}$ satisfying Assumptions 1.iii. and 1.iv., we have $\frac{(1-c_{{\nu}}){\nu}_{0}^{2}-{\nu}_{i}^{2}}{2(1-c_{{\nu}})}>0$ and $\log\frac{c_{{\nu}}{\nu}_{i}^{2}}{(c_{L}^{2}+c_{{\nu}}c_{L}){\nu}_{0}^{2}}-\frac{{\nu}_{i}^{2}}{2(1-c_{{\nu}}){\nu}_{0}^{2}}+\frac{(1-c_{{\nu}}){\nu}_{0}^{2}}{2{\nu}_{i}^{2}}>0$ are both positive constants.
@@
where ${\mathbf{R}}\in{\mathbb{R}}^{d\times r}$ an orthonormal basis of the vector space $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$ and ${\mathbf{N}}\in{\mathbb{R}}^{d\times n}$ an orthonormal basis of the null space of $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$. Then, conditioned on $\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\left(\frac{3{\nu}_{0}^{2}}{4}+\frac{{\nu}_{\max}^{2}}{4(1-c_{{\nu}})}\right)d$, we prove that $\left\|{\mathbf{n}_{t}}\right\|$ remains large with high probability.
@@
To show $\frac{\left\|{\nabla_{\mathbf{x}}P^{(i)}({\mathbf{x}})}\right\|}{P({\mathbf{x}})}\leq\exp(-\Omega(d))$, from Assumptions 1.ii. and 1.iii. we have
@@
Armed with the above lemmas, we are now ready to establish Theorem 5 by induction. Please note that we recycle some lemmas from the proof of Theorem 2 by substituting ${\nu}_{\max}^{2}$ with $\frac{{\nu}_{\max}^{2}}{1-c_{{\nu}}}$. Suppose the theorem holds for all $T$ values of $11,\cdots,T-$ 1. We consider the following 3 cases:
@@
Therefore we complete the proof of Theorem 5. ∎
B.2 Proof of Theorem 6